### v4

# Vivid Emotions

Abstract
GPT-4o, an omni-modal model that enables vocal conversations with diverse emotions and tones, marks a milestone for omni-modal foundation models. However, empowering Large Language Models to perceive and generate images, texts, and speeches end-to-end with publicly available data remains challenging for the open-source community. Existing vision-language models rely on external tools for speech processing, while speech-language models still suffer from limited or totally without vision-understanding capabilities. To address this gap, we propose the EMOVA (EMotionally Omni-present Voice Assistant), to enable Large Language Models with end-to-end speech abilities while maintaining the leading vision-language performance. With a semantic-acoustic disentangled speech tokenizer, we surprisingly notice that omni-modal alignment can further enhance vision-language and speech abilities compared with the bi-modal aligned counterparts. Moreover, a lightweight style module is introduced for the flexible speech style controls including emotions and pitches. For the first time, EMOVA achieves state-of-the-art performance on both the vision-language and speech benchmarks, and meanwhile, supporting omni-modal spoken dialogue with vivid emotions.
00footnotetext: Equal contribution, listed in alphabetical order by surname.00footnotetext: Corresponding authors: {honglanqing,houlu3,xu.hang}@ huawei.com
1 Introduction
OpenAI GPT-4o [70], a novel milestone for the omni-modal foundation models, has rekindled people’s attention on intelligent assistants that can see (i.e., perceiving fine-grained visual inputs), hear (i.e., understanding vocal instructions) and speak (i.e., generating vocal responses) simultaneously. Existing Multi-modal Large Language Models (MLLMs) mostly focus on two modalities only, either vision-language [1, 40] or speech-language [11, 87], demonstrating severe demands for omni-modal models with visual, language and speech abilities. How to empower Large Language Models (LLMs) to effectively process omni-modal data in an end-to-end manner remains an open question.
Figure 1: EMOVA is the very first omni-modal LLM with state-of-the-art performance on both vision-language and speech benchmarks simultaneously. See detailed results in Table 2.
Existing omni-modal LLMs [10, 19] generally build on Vision LLMs and integrate the speech modality by adopting a speech encoder like Whisper [74], which extracts continuous speech features, similar to how images are processed, to enable speech understanding. These models, however, still rely on external Text-to-Speech (TTS) tools for generating speech responses, limiting their ability to support real-time interactions. AnyGPT [93], on the contrary, opts for a fully discretization way, which first discretizes all data modalities (i.e., images, texts, and speeches), followed by omni-modal auto-regressive modeling. This enables AnyGPT to handle multiple modalities with a unified end-to-end framework, facilitating real-time interactions with the help of streaming decoding. However, the discrete vision tokenizer used by AnyGPT struggles to capture visual details, especially for high-resolution images, falling far behind its continuous counterparts on vision-language benchmarks. Furthermore, none of the existing works explore the speech style controls (e.g., emotions and pitches) with LLMs. Thus, our question arises: How to build an end-to-end omni-modal LLM enabling emotional spoken dialogue while maintaining state-of-the-art vision-language performance?
In this paper, we propose EMOVA (EMotionally Omni-present Voice Assistant), a novel end-to-end omni-modal LLM with state-of-the-art vision-language and speech capabilities while supporting emotional spoken dialogue. Fig. 2 shows an overview of the model framework. A continuous vision encoder captures fine-grained visual details, while a discrete speech tokenizer and detokenizer empower end-to-end speech understanding and generation. Specifically, the speech-to-unit (S2U) tokenizer tokenizes the input speech waveforms to discrete speech units, while the unit-to-speech (U2S) detokenizer reconstructs the speech waveforms from LLM’s output speech units. To seamlessly integrate speech modality with LLMs, we meticulously design a semantic- acoustic disentangled speech tokenizer to decouple the semantic contents and acoustic styles of input speeches [77], where 1) semantic content (i.e., what it says) captures the semantic meanings of the input speeches, which is finally discretized and aligned with LLMs, while 2) acoustic style (i.e., how it says) captures the diverse speech styles (e.g., emotions and pitches). Utilizing the semantic-acoustic disentanglement of our speech tokenizer, we further introduce a lightweight style module to support spoken dialogue with vivid emotions and pitches. As in Sec. 4.1, this disentanglement design better facilitates the modality alignment among texts and speeches while maintaining flexibility for diverse speech style controllability and personalization.
With EMOVA’s end-to-end omni-modal framework, we empirically demonstrate publicly available bi-modal image-text and speech-text data are sufficient for the omni-modal alignment with the text modality as a bridge, eliminating the need for omni-modal data (i.e., image-text-speech), which is usually scarce. Surprisingly, we find that the omni-modal alignment can further improve both the vision-language and speech capabilities via joint optimization, even when compared with their bi-modal aligned counterparts. Ultimately, only a small amount of omni-modality samples are required to teach the model to respond in the desired format. For the first time, EMOVA obtains state-of-the-art results on both vision-language and speech benchmarks (see Table 2). The main contributions of this work contain three parts:
  1. 1.
    We propose EMOVA, a novel end-to-end omni-modal LLM that can see, hear and speak. A continuous vision encoder with a semantic-acoustic disentangled speech tokenizer is adopted for seamless omni-modal alignment and diverse speech style controllability.
  2. 2.
    We introduce an efficient text-centric omni-modal alignment which can further enhance the vision-language and speech abilities, surpassing their bi-modal aligned counterparts (i.e., image-text only and speech-text only).
  3. 3.
    For the first time, EMOVA obtains state-of-the-art comparable results on both the vision-language and speech benchmarks simultaneously and further supports flexible spoken dialogues with vivid emotions.
Method Visual   Text Speech
Understand Gen. Emotion
Vision Large Language Models
LLaVA [49]
Intern-VL [10]
Speech Large Language Models
Qwen-Audio [11]
Mini-Omni [87]
LLaMA-Omni [17]
Omni-modal Large Language Models
VITA [19, 20]
Ola [55]
Any-GPT [93]
Baichuan-Omni [43]
EMOVA (ours)
Table 1: Comparison of Multi-modal Large Language Models. EMOVA is the very first Omni-modal LLM capable of emotional spoken dialogue with state-of-the-art vision-language and speech capabilities simultaneously. “Gen.” stands for Generation.
2 Related Work
Vision Large Language Models
(VLLMs) integrate the vision modality into LLMs [81, 7], enabling the advanced understanding and reasoning over visual instructions [49, 1, 26, 27]. Recent VLLM works can be categorized into three directions, 1) Vision encoders [71, 5, 6] are enhanced and aggregated for robust representations [46, 45, 80]. 2) High-resolution methods are proposed to overcome the fixed resolution of pre-trained vision encoders (e.g., $336\times 336$ for CLIP [73]), enabling LLMs to perceive fine-grained visual information [48, 14, 31, 58]. 3) High-quality instruction data is essential for VLLMs to generate accurate and well-formed responses [36, 40, 10]. Besides achieving state-of-the-art vision-language performance, we further introduce speech understanding and generating abilities to EMOVA.
Figure 2: Model architecture of EMOVA. The vision encoder extracts continuous visual features, which are projected into the textual embedding space as visual tokens, while the input speech is encoded and quantized into discrete speech units. Given the omni-modal inputs, EMOVA can generate both textual and speech responses with vivid emotional controls. Check Sec. 3 for more architectural details.
Speech Large Language Models
(SLLMs) empower the speech interaction with LLMs. Continuous SLLMs [84, 11] adopt the speech encoders [74] to extract continuous speech embeddings for LLM, which, however, only support speech understanding, relying on external TTS modules for speech generation, and therefore, hampering real-time interaction. Discrete SLLMs [95], instead, first discretize speech signals with speech tokenizers, followed by auto-regressive modeling. Recent works [17, 87] further combine the continuous speech encoders with discrete speech tokenizers for better results. Although effective, none of the existing works explore speech style controllability in SLLMs (e.g., emotions and pitches), which is essential for spoken dialogue.
Omni-modal Large Language Models
support visual, text, and speech abilities with a unified architecture. Similar to continuous SLLMs, InternOmni [10] and VITA [19] connect a speech encoder with VLLMs, supporting speech understanding only. Instead, AnyGPT [93] proposes a unified architecture to discretize and conduct auto-regressive modeling for image, text, and audio simultaneously, which, however, suffers from inevitable information loss brought by discretization, especially for the high-resolution visual inputs. Our EMOVA is the very first unified Omni-modal LLM with state-of-the-art vision-language and speech performance at the same time.
3 Architecture
3.1 Formulation
Denote LLM as $f(\cdot)$ and text, visual and speech inputs as $\textbf{X}_{T}$, $\textbf{X}_{V}$ and $\textbf{X}_{S}$, respectively. $\textbf{X}_{T}$ is converted to discrete tokens $\textbf{U}_{T}$ via a text tokenizer [21], while the $\textbf{X}_{V}$ is first encoded with a vision encoder $v(\cdot)$ as $\textbf{E}_{V}=v(\textbf{X}_{V})$, and then projected into the textual embedding space with a projector $p(\cdot)$ as $\textbf{H}_{V}=p(\textbf{E}_{V})$. As for the speech input $\textbf{X}_{S}$, a Speech-to-Unit (S2U) procedure is required. Specifically, $\textbf{X}_{S}$ first goes through a speech encoder $s(\cdot)$ as $\textbf{E}_{S}=s(\textbf{X}_{S})$, which is then discretized by the quantizer $q(\cdot)$ as $\textbf{U}_{S}=q(\textbf{E}_{S})$. The LLM $f(\cdot)$ is then trained to compute the joint probability of the output texts $\textbf{U}_{T}^{o}$ and speech units $\textbf{U}_{S}^{o}$ as
$\mathbb{P}(\textbf{U}_{T}^{o},\textbf{U}_{S}^{o}|\textbf{U}_{omni})=\prod_{i=1}^{L}\mathbb{P}(\boldsymbol{x}_{i}|\textbf{U}_{T,<i}^{o},\textbf{U}_{S,<i}^{o},\textbf{U}_{omni}),$ (1)
where $\boldsymbol{x}_{i}\in\textbf{U}_{T}^{o}\cup\textbf{U}_{S}^{o}$, $L=|\textbf{U}_{T}^{o}|+|\textbf{U}_{S}^{o}|$ and $\textbf{U}_{omni}=\textbf{U}_{T}\cup\textbf{U}_{S}\cup\textbf{H}_{V}$, which stands for the omni-modal inputs. The output response units $\textbf{U}_{S}^{o}$ are then recovered into the output speech waveform $\textbf{Y}_{S}^{o}$ via a Unit-to-Speech (U2S) decoder $d(\cdot,\cdot)$ with an emotion style embedding $\textbf{E}_{style}^{o}$ to realize the vivid emotional spoken dialogue controllability (Sec. 3.2).
LLM.
We utilize the Qwen-2.5 [79] model families as the base LLMs of EMOVA with three configurations (i.e., 3B, 7B, and 72B) for usage under different budgets.
Vision encoder and projector.
We use the QwenViT [78] as the visual encoder $v(\cdot)$ with an MLP vision projector $p(\cdot)$ with a 4×4$\times$4 × downsample rate for all variants of EMOVA.
[…]
3.2 Speech Tokenization
Speech-to-unit (S2U) tokenizer.
Following [77], we use the SPIRAL [32] architecture for the speech encoder $s(\cdot)$ to capture both phonetic and tonal information, which is then discretized by the quantizer $q(\cdot)$ with finite scalar quantization (FSQ) [64]. The size of the speech codebook is 4,096, while the sample rate is 25 tokens per second. Once discretized, the speech modality can be integrated into LLMs by concatenating the text vocabulary and speech codebook.
Our S2U tokenizer provides the following advantages: 1) Data efficiency: after pre-training on large-scale unlabeled speech data, it requires only a small amount of speech-text pair data for easy adaptation. 2) Bilingual: the speech codebook is shared among languages (i.e., English and Chinese), sharing unit modeling capabilities across languages. Check more training details and comparisons in Appendix A.1.
Semantic-acoustic disentanglement.
To align the speech units seamlessly with the highly semantic embedding space of LLMs, we opt for decoupling the semantic contents and acoustic styles of input speeches. Given input speechs $\textbf{X}_{S}$, both semantic embedding $\textbf{E}_{semantic}$ and style embeddings $\textbf{E}_{style}$ are extracted separately as
$\{\textbf{E}_{semantic},\textbf{E}_{style}\}=\textbf{E}_{S}=s(\textbf{X}_{S}).$ (2)
Only $\textbf{E}_{semantic}$ is quantified by $q(\cdot)$ to generate the speech units $\textbf{U}_{S}$. By using different $\textbf{E}_{style}$ while maintaining the same $\textbf{E}_{semantic}$, we can easily control the recovered speech styles without disturbing the semantic contents of recovered speeches. Moreover, the disentanglement facilitates modality alignment among speeches and texts, as later in Sec. 4.1.
Unit-to-speech (U2S) detokenizer with style controls.
Building on VITS [35], our U2S detokenizer adopts a conditional VAE architecture (cf., Fig. 7). To achieve flexible style controls, we utilize the semantic-acoustic disentanglement of our S2U tokenizer (as discussed above) and adopt a novel style embedding to control the speech styles (e.g., genders, emotions, and pitches). Specifically, the LLM $f(\cdot)$ is trained to generate both the output speech units $\textbf{U}_{S}^{o}$ and a style label. The speech units $\textbf{U}_{S}^{o}$ are converted to the unit embeddings $\textbf{E}_{semantic}^{o}$, while the style label is utilized to generate a unique style prototype $\textbf{E}_{style}^{o}$. Both $\textbf{E}_{semantic}^{o}$ and $\textbf{E}_{style}^{o}$ are taken as inputs to speech decoder $d(\cdot,\cdot)$ to synthesize output speeches $\textbf{Y}_{S}^{o}=d(\textbf{E}_{semantic}^{o},,\textbf{E}_{style}^{o})$.
Our U2S detokenizer is pre-trained on LibriTTS [92] and AISHELL-1 [2] and subsequently fine-tuned on synthetic style-rich speech data. Due to the scarcity of real-life style-rich data, we utilize TTS tools [16] to synthesize the speech samples diverse in genders, pitches, and emotions. As for the style prototypes, Emotion2Vec [59] is adopted to select the most representative samples with the highest confidence in conveying the desired style. Our empirical results reveal that even one representative style reference speech has been sufficient to control the speech styles flexibly and precisely. Check Appendix A.2 for more details.
Figure 4: Demonstration of EMOVA omni-modal instruction tuning. 1) To support emotional spoken dialogues, EMOVA is trained to explicitly select speech style labels with output speech units. 2) For ease of parsing, the data elements are organized in the JSON format.
4 Training Omni-modal LLMs
To achieve omni-model alignment, it is ideal to use large-scale omni-modal image-text-speech data, which, however, is either without reach due to copyrights [67] or limited in quality [65]. An alternative is to use the existing image-text data with the TTS-synthesized speeches, which is not only computationally expensive but also hampers data diversity, as most TTS tools generate speeches in similar patterns. Recent works [10, 19] choose to integrate the speech modality into a well-structured VLLM via a sequential training manner with bi-modal alignment data. However, the relationships among modalities and how to effectively leverage multiple bi-modal alignment datasets remain unclear.
In this work, we explore omni-modal text-centric alignment by utilizing the publicly available bi-modal alignment datasets, including image-text (e.g., image captioning) and speech-text (e.g., ASR and TTS) datasets. With text modality as a bridge, our EMOVA ultimately becomes a unified system capable of understanding and generating multiple modalities in a coherent and integrated manner. In Sec. 4.1, we first explore the following three questions:
  1. 1.
    Does the integration of the speech modality conflict with the vision-language capabilities?
  2. 2.
    Is sequential alignment of multiple modalities optimal?
  3. 3.
    How to represent speech modality to foster omni-modal alignment?
We then discuss the omni-modal instruction tuning pipeline and the overall training paradigm of EMOVA in Sec. 4.2 and Sec. 4.3, respectively.
4.1 Omni-modal Text-Centric Alignment
Settings.
[…]
Evaluation.
For speech abilities, we evaluate the aligned model’s performance on the ASR task of LibriSpeech [72], while for the vision-language, we fine-tune the model with a small amount of high-quality visual instruction data (i.e., 665K SFT data from the ShareGPT4V [9]) and evaluate the fine-tuned model on common vision-language benchmarks. Check Appendix C for evaluation details. Fig. 3 shows the comparison among different paradigms on vision-language (left and middle) and ASR (right, where we report the $1-\text{WER}$ value for better readability) benchmarks, from which we can derive the following observations:
Observation 1: image-text and speech-unit-text data benefit each other.
[…]
Observation 2: semantic-acoustic disentanglement benefits omni-modal alignment.
We find that 1) Joint outperforms Joint-entangled on vision-language benchmarks, and 2) in the speech tasks, Joint maintains significant advantages over its entangled counterpart. This can be attributed to the semantic-acoustic disentanglement which makes speech units more analogous to languages.
Benchmarks EMOVA EMOVA EMOVA Gemini GPT- GPT- Whisper Mini- VITA VITA Baichuan-
3B 7B 72B Pro 1.5 4V 4o Large Omni2 8x7B 1.5 Omni-7B
MME 2175 2317 2402 - 1927 2310 - - 2097 2311 2187
MMBench 79.2 83.0 86.4 - 75.0 83.4 - - 71.8 76.6 76.2
SEED-Image 74.9 75.5 76.6 - 71.6 77.1 - - 72.6 74.2 74.1
MM-Vet 57.3 59.4 64.8 - 67.7 - - - 41.6 51.1 65.4
RealWorldQA 62.6 67.5 71.0 68.7 61.4 75.4 - - 59.0 66.8 62.6
TextVQA 77.2 78.0 81.4 73.5 77.4 - - - 71.8 74.9 74.3
ChartQA 81.5 84.9 88.7 81.3 78.5 85.7 - - 76.6 79.6 79.6
DocVQA (test) 93.5 94.2 95.9 86.5 88.4 92.8 - - - - -
InfoVQA (test) 71.2 75.1 83.2 72.7 - - - - - - -
OCRBench 803 814 843 - 656 736 - - 678 752 700
AI2D 78.6 81.7 85.8 80.3 78.2 84.6 - - 73.1 79.3 -
ScienceQA-Img 92.7 96.4 98.2 - 75.7 - - - - - -
MMMU 45.8 49.8 59.7 58.5 56.8 69.2 - - 47.3 52.1 47.3
MathVista 62.6 65.5 69.9 52.1 49.9 63.8 - - 44.9 66.2 51.9
Mathverse 31.4 40.9 50.0 - 33.6 - - - - - -
[…]
Table 2: Comparison on vision-language and speech benchmarks. 1) EMOVA outperforms GPT-4o/4V and Gemini Pro 1.5 on 11 of the 15 vision-language benchmarks, providing a powerful open-sourced alternative. 2) Meanwhile, our EMOVA achieves state-of-the-art performance on Librispeech, surpassing its speech and omni-modal counterparts significantly. : reported by [43].
Observation 3: sequential alignment is not optimal.
[…]
4.2 Omni-modal Instruction Tuning
After the omni-modal text-centric alignment in Sec. 4.1, the model learns the fundamental vision-language (e.g., captioning) and speech capabilities (e.g., ASR and TTS). However, instruction tuning is essential to better follow complicated user instructions and respond with vivid emotions.
Emotion-enriched instruction data synthesis.
[…]
Data organization and the chain of modality.
The omni-modal instruction data can be represented as $D_{\text{omni}}=\{(x_{V},u_{S},x^{o}_{T},c^{o}_{\text{style}},u^{o}_{S})_{i}\}_{i=1}^{N}$, where the input consists of the optional queried image $x_{V}$ and the speech units of the instruction $u_{S}$, while the output consists of the textual response $x^{o}_{T}$, the predicted speech style labels $c^{o}_{\text{style}}$, and the output speech unit $u^{o}_{S}$. Note that we train EMOVA to explicitly select styles (e.g., emotions and pitches), which are utilized to determine the corresponding style embedding for the U2S detokenizer (Sec. 3.2). Furthermore, since directly generating speech responses is challenging, we decompose the speech response procedure into three primary steps: 1) recognizing user instructions into texts; 2) generating textual responses based on the recognized instructions; 3) generating the style labels and response speech units based on the textual responses. For ease of parsing during deployment, the target outputs are formatted as JSON, as in Fig. 4.
Datasets End-$\uparrow$ Text response Style Categorization Recognition/Synthesis
[…]
Speech-Image-EN 7.45 7.56 7.95 82.50 97.70 2.40 3.20
Speech-Text-EN 6.85 6.90 7.38 81.20 84.70 6.90 2.90
Speech-Image-ZH 6.48 7.02 6.82 77.60 95.90 1.70 12.00
Speech-Text-ZH 5.25 5.58 6.60 80.90 93.20 10.70 12.20
Table 3: Evaluation of EMOVA-7B on Speech Dialogue. By default, we evaluate on the corresponding test set of the evaluated datasets.
4.3 Overall Training Paradigm
Inspired by [9], a three-stage training paradigm is adopted,
  • Stage-1: Vision-language pre-alignment. The purpose is to align the visual features into the embedding space of LLMs. Only the vision projector $p(\cdot)$ is trained.
  • Stage-2: Omni-modal text-centric alignment. This stage performs vision-language and speech-language alignment jointly. We train the LLM $f(\cdot)$, vision projector $p(\cdot)$, and the deeper half of vision encoder $v(\cdot)$ layers.
  • Stage-3: Omni-modal instruction tuning. We organize different datasets with various types of instructions to learn generalization across tasks, as detailed in Sec. 5.1.
5 Experiments
5.1 Training configuration
Stage-1.
In this stage, we only train the parameters of the vision projector $p(\cdot)$ for vision-language pre-alignment with the LCS-558K dataset [49], with the high-resolution image-slicing strategy [48] adopted.
Stage-2.
We assemble a unified dataset with 7.4M samples for both the image-text and speech-text alignment, as summarized in Fig. 9. Specifically, we utilize pre-training datasets from ShareGPT4V [9], ALLaVA [4] (both the original English version and the Chinese version translated on our own), and ShareGPT-4o [12] for general perception, while for the OCR capabilities, we leverage SynthDog [34], MMC-Alignment [47], K12 Printing, and the UReader Text Reading subset [89]. Moreover, we adopt the 2,000 hours of ASR and TTS data from LibriSpeech [72] and AISHELL-2 [15] for speech-text alignment, and to preserve the language capabilities of LLMs, we further incorporate the text-only data from Magpie Pro [88]. Check more details in Fig. 9.
Stage-3.
We collect the EMOVA-SFT dataset consisting of 4.4M multi-task omni-modal samples (see Fig. 8). We start by gathering high-quality open-sourced visual instruction datasets, including ShareGPT4V [9], InternVL [10], Meteor [38], Idefics-2 [36], Cambrian [80], and LLaVA-Onevision [40], followed by quality checking, re-formatting all data samples with a unified template, and removing the duplicated data. For speech, we include the training split of EMOVA omni-model instruction data (cf., Sec. 4.2), with 10% of speech alignment datasets to maintain ASR and TTS performance. We train with 128 Ascend 910B (64GB) NPUs in parallel (check more details in Table 5).
5.2 Comparison to SOTA Models
Experimental results are provided in Table 2. We compare a wide range of state-of-the-art VLLMs, including Gemini Pro 1.5 [75], GPT-4V [69], GPT-4o [70], together with the Speech LLM (i.e., Mini-Omni2 [87]) together with the ASR expert Whisper-Large [74], and the omni-modal LLMs (i.e., VITA-8x7B [19], VITA-1.5 [20] and Baichuan-Omni [43]).
Comparison with SOTA VLLMs.
As an omni-modal model, EMOVA obtains comparable performance with the state-of-the-art VLLMs on multiple vision-language benchmarks, while showing superior proficiency in solving math problems needing precise visual content interpretation. Our EMOVA-7B surpasses GPT-4V by +7.3 on MathVerse, and our EMOVA-72B exceeds GPT-4o by +6.1 on MathVista. On 11 out of the 15 benchmarks, our EMOVA outperforms both GPT-4o/4V and Gemini Pro 1.5, providing a powerful open-sourced alternative.
Comparison with SOTA omni-modal LLMs.
Under the similar 7B capacity, EMOVA-7B is 220 points higher than VITA on MME, surpassing VITA by 20.6% on OCRBench (814 vs. 678). Moreover, it surpasses Baichuan-Omni-7B, a more recent Omni-modal LLM, on nearly all the evaluated benchmarks, underscoring the effectiveness of our approach and potential to push the boundary of omni-modal abilities.
Moreover, ASR results show that EMOVA significantly outperforms the most recent omni-modal model VITA, even surpassing its SLLM counterparts Mini-Omni2 and ASR-specialized Whisper-Large, showing the effectiveness of the semantic-acoustic disentanglement and omni-modal mutual benefits. We further report TTS results in Table 6, and provide the qualitative examples in Fig. 5 and Appendix G. For the first time, EMOVA obtains state-of-the-art results on both the speech and vision-language benchmarks.
Figure 5: EMOVA engages in omni-modal emotional spoken dialogue expressing sadness.
5.3 Evaluation of Emotion-Rich Spoken Dialogue
In this section, we evaluate the end-to-end spoken dialogue abilities of EMOVA-7B. As stated in Sec. 4.2, the model takes an input image $x_{V}$ and user instructions in the form of speech units $u_{S}$, and outputs text responses, style labels, and corresponding speech units. To ensure comprehensive evaluation, we propose the following evaluation metrics:
  1. 1.
    End-to-end spoken dialogue score assesses the model’s dialogue performance based on the generated speeches, with a score ranging from 0 to 10, reporting the average.
  2. 2.
    Unit-input-text-output score focuses on the quality of the textual responses of LLM when the inputs are speech units, bypassing errors from speech synthesis.
  3. 3.
    Text-input-text-output score inputs the ground-truth user instruction texts and evaluates the model’s text outputs. This helps disentangle the impact of speech recognition errors and eliminates the effect of JSON format.
  4. 4.
    ASR and TTS evaluate how accurately the model recognizes the speech units and how effectively it generates speech units from text. See Appendix D for more details.
  5. 5.
    Style label classification accuracy evaluates the accuracy in selecting the appropriate speech style labels.
  6. 6.
    Style controllablity assesses the controllability of U2S detokenizer with the given conditional style labels using the confusion matrix comparing the generated and recognized style labels. See Appendix D for more details.
Due to the lack of emotionally rich spoken dialogue evaluation datasets, we split a test set from our synthesized omni-modal instruction-tuning data (Sec. 4.1). GPT-4o is used for automated evaluation. See details in Appendix D.
Results.
Table 3 the spoken dialogue performance.
(i) By comparing the end-to-end dialogue score with the unit-input-text-output score, we notice that the two scores are closely aligned, with a maximum gap of only $0.33$, except for Speech-Image-ZH. TTS-WER/CER is generally low for English, revealing that EMOVA can synthesize accurate speech based on textual responses, which, however, is harder for Chinese, which we attribute to its complexity. It includes tasks such as generating poetries and answering riddles, resulting in more intricate responses.
Figure 6: Confusion matrix between the generated and recognized emotions. The emotions generated by our U2S detokenizer are recognized with high probability. Best view with zooming in.
(ii) Comparing the unit-input-text-output score with the text-input-text-output score, we notice that their differences correlate with the ASR performance of the speech instructions, especially for Speech-Text-EN and Speech-Text-ZH, which involve more complex instructions.
Our EMOVA-7B reports inferior ASR performance ($6.9$ and $10.7$, respectively) compared to other datasets ($2.4$ and $1.7$). Consequently, when we replace speech instructions with ground-truth transcriptions, EMOVA shows significant improvements from unit-input to text input score. On the contrary, for datasets with accurate ASR performance, the results are quite similar, suggesting EMOVA retains robust dialogue capabilities when using the JSON format.
(iii) Examining the classification accuracy of style labels, we find that EMOVA performs satisfactorily in classifying emotions and pitches during speech conversations, achieving an accuracy of over 75%. The confusion matrix comparing the conditional and recognized emotion labels is shown in Fig. 6. The results indicate that the four emotions are recognized with high probabilities, with three achieving over 80% accuracy. This demonstrates that our U2S detokenizer effectively controls common emotions, endowing the synthesized speech with vivid emotional expression.
6 Conclusion
Our work builds EMOVA, a novel end-to-end omni-modal large language model that effectively aligns vision, speech, and text simultaneously. With text as a bridge, we show that omni-modal alignment is achievable without relying on omni-modal image-text-speech data, meanwhile, enhancing both vision-language and speech abilities. For the first time, EMOVA achieves state-of-the-art performance on both vision-language and speech benchmarks, setting a new standard for versatile omni-modal interactions.
Acknowledgments.
We gratefully acknowledge supports of MindSpore, CANN (Compute Architecture for Neural Networks) and Ascend AI Processor used for this research. This work has been made possible by a Research Impact Fund project (RIF R6003-21) and a General Research Fund project (GRF 16203224) funded by the Research Grants Council (RGC) of the Hong Kong Government. This work was partially supported by the Research Grants Council of the Hong Kong Special Administrative Region (Grants C7004-22G-1 and 16202523) and the Joint Centre for Artificial Intelligence (Grant FB453). This work is supported by National Key Research and Development Program of China (2024YFE0203100) and National Natural Science Foundation of China (No. 62441615, 62201484 and 62136005).
Appendix
Appendix A More on Speech Tokenizer
A.1 Speech-to-Unit (S2U) Tokenizer
Overview.
To process the speech input $\textbf{X}_{S}$, our S2U tokenizer consists of a speech encoder $s(\cdot)$ with a quantization module $q(\cdot)$. First, the speech input is passed through $s(\cdot)$, producing a continuous latent representation $\textbf{E}_{S}=s(\textbf{X}_{S})$. Then, the quantization module $q(\cdot)$ converts $\textbf{E}_{s}$ into discrete units $\textbf{U}_{S}=q(\textbf{E}_{s})$. The final output is an ID sequence $\textbf{U}_{S}=[u_{1},u_{2},\cdots]$, where each $u_{i}$ corresponds to a unique speech unit in the speech codebook.
After this S2U extraction procedure, the speech is represented by quantized IDs instead of quantized embeddings. For example, a speech signal is represented as an ID sequence like $[782,463,550,\cdots]$, which can be treated as a special form of text. As a result, to integrate speech tokens into LLM $f(\cdot)$, we only need to expand the LLM’s original vocabulary $V_{T}$ by adding a set of speech unit tokens $V_{S}$, similar to [95]. The expanded vocabulary is thus the union $V=V_{T}\cup V_{S}$. In this work, the same codebook is shared across multiple languages, such as English and Chinese, enabling the unit modeling abilities to be shared across languages. This design enables simply integration of the speech modality into the LLM with minimal data requirements (see experiments in Sec. 4.1).
Training of S2U.
The training of the S2U encoder involves three primary components: the speech encoder, the quantization module, and the phone decoder. First, the speech encoder is trained with a large amount of unlabeled speech with contrastive loss in a self-supervised learning manner [77]. The dataset utilized is the 10000-hour English speeches from GigaSpeech [3] and the 10000-hour Chinese speeches from the WenetSpeech [94], both of which encode large variations in speakers, styles, and acoustic environments. Subsequently, the speech encoder, along with the quantization module and phone decoder, are optimized on a small amount of speech-text pair data, which is derived from the train-clean-100 subset of LibriSpeech [72] in English and AISHELL-1 [2] in Chinese. The phone label is obtained using the phone set in Charsiu [99]. During training, the speech encoder encodes input speeches into continuous latent representations that are rich in both phonetic and tonal information. Then, the quantization module is used to convert the continuous outputs from the speech encoder into discrete units. Finally, the phone decoder converts the quantized embeddings into a sequence of non-tonal/tonal phones, ensuring that the speech units capture necessary information related to semantic contents in both non-tonal and tonal languages. After training, only the speech encoder and the quantization module are used in EMOVA.
Comparisons with SpeechTokenizer in AnyGPT.
Our S2U tokenizer differs from the SpeechTokenizer [97] adopted in AnyGPT [93], in the following aspects:
(1) SpeechTokenizer encodes both semantic contents and acoustic details of speeches, while our S2U tokenizer focuses solely on semantic contents. This design reduces the LLMs’ burden of disentangling different aspects of speech information, facilitating the modality alignment between speech and text modalities during LLM training.
(2) Compared with SpeechTokenizer, our S2U tokenizer offers a more concise representation and helps to simplify and accelerate generation. SpeechTokenizer employs tokens from eight RVQ layers with a 50Hz frame rate to represent speech, thus a 10-second speech corresponds to 500 × 8 = 4000 tokens. However, we reduce the frame rate from 50Hz to 25Hz and utilize only one token to represent each frame, and thus, a 10-second speech can be represented by only 250 tokens. Moreover, AnyGPT requires a two-stage generation process, involving autoregressive (AR) semantic token generation followed by the non-autoregressive (NAR) acoustic token generation. Instead, we only need to generate speech units capturing the semantic contents in a fully AR manner.
(3) SpeechTokenizer lacks an explicit structure design to deal with tonal languages like Chinese, therefore, the processing ability in Chinese is not demonstrated in either SpeechTokenizer or AnyGPT. In contrast, our S2U tokenizer incorporates training constraints to better capture tone variation in phone, making it effective for both the non-tonal and tonal languages. This further enhances EMOVA’s multilingual speech processing capabilities, enabling it to effectively handle both English and Chinese.
In summary, our S2U tokenizer improves the compactness and generality of speech representation, facilitates LLM training, and enhances its multilingual speech ability. Experimental results show that our model significantly outperforms AnyGPT in ASR tasks, as shown in Table 6.
Figure 7: U2S detokenizer with style control.
A.2 Unit-to-Speech (U2S) Detokenizer with Style Control
Overview.
The LLM, along with the vision encoder and speech tokenizer, is trained end-to-end to generate responses in the form of the speech units, given the input images and speeches. Specifically, the output speech units can be obtained via $\textbf{U}_{S}^{o}=f(\textbf{U}_{T},\textbf{U}_{S},\textbf{H}_{V})$, followed by a U2S detokenizer to convert the discrete speech units $\textbf{U}_{S}^{o}$ into the final output speech waveforms.
The proposed U2S detokenizer involves three core modules: the speech unit encoder $e(\cdot)$, the speech style encoder $g(\cdot)$, and the speech decoder $d(\cdot,\cdot)$. First, the speech unit encoder converts the speech units $\textbf{U}_{S}^{o}$ into unit embeddings $\textbf{E}_{unit}^{o}$. Meanwhile, the style encoder $g(\cdot)$, adopting the structure of Meta-StyleSpeech [66], is utilized to extract a style embedding $\textbf{E}_{style}^{o}$ from the chosen reference speech. Lastly, the speech decoder $d(\cdot,\cdot)$ reconstructs the speech waveform $\textbf{Y}_{S}^{o}$ from the unit embedding $\textbf{E}_{unit}^{o}$ and style embedding $\textbf{E}_{style}^{o}$.
Training of U2S.
Training a U2S detokenizer with emotion controls is challenging considering the lack of labeled emotional speech data since most open-source speech data is predominantly neutral in emotion or lacks emotion labels. Due to the limited availability of emotion-rich data, we utilize TTS tools [16] to generate a small set of style-rich speech samples diverse in speaker identities, genders, emotions, and pitches. Our U2S detokenizer is first pre-trained on LibriTTS [92] and AISHELL-1 [2] to acquire fundamental speech synthesis capabilities, and subsequently, the synthesized style-rich speech data is utilized to fine-tune the U2S detokenizer, enhancing its controllability over diverse speech styles.
Style Prototypes.
To better facilitate controls of genders, emotions, and pitches, inspired by [66] that a small number of style reference speeches can effectively transfer the target styles, we adopt a “store-for-usage" manner, i.e., we construct a style prototype codebook in advance for speech style assignation. Specifically, we synthesize $K$ reference candidates with external TTS tools for each possible combination of the following styles: two genders (male, female), four emotions (neutral, happy, sad, angry), and three pitches (normal, high, low), leading to 24 unique styles and $24\times K$ candidates. Empirically we find that genders and pitches are easy to control using any of the candidate references, while the emotion intensity varies across speeches. To tackle this, we adopt Emotion2Vec [59], a powerful speech emotion recognition (SER) tool, to measure the emotion intensity of each candidate reference, and rank them in terms of the confidence of the desired emotion. We select the Top-1 candidate reference in each combination style to be the prototype of this condition. Finally, the most representative 24 reference speeches are selected from the $24\times K$ candidates.
Appendix B More on Omni-modality
B.1 Omni-modal Text-Centric Alignment
Modality alignment data
is summarized in Fig. 9.
Experiments on Omni-modal Alignment Paradigms.
The training configuration adopted in Sec. 4.1 is mostly identical to Table 5 except we use a unique resolution of 448 for all stages and replace EMOVA-SFT-4M in Stage-3 with ShareGPT4V [9] for efficiency.
Given the space constraints, the evaluation benchmarks in Fig. 3 represent selected benchmarks from each category in Table 2. Specifically, for general image perception and understanding, we choose MMBench and MM-Vet; for mathematical problem solving, we adopt MathVista (testmini); for science understanding, we select ScienceQA-Img; and for automatic speech recognition (ASR), we utilize the test-clean split of the LibriSpeech dataset.
B.2 Omni-modal Instruction Data Synthesis
Dataset construction.
To obtain emotion and pitch labels, we leverage GPT-4o using the prompt in Fig. 23. Table 4 shows the distribution of speech styles of our speech instruction dataset.
Figure 8: Overview of the data composition for EMOVA omni-modal instruction tuning. (Left) Distribution of instruction data across categories, with the outer circle representing overall categories and the inner circle depicting subset distributions. (Right) Quantitative breakdown of data sources.
Figure 9: Overview of EMOVA omni-modal alignment data composition.
Detailed data organization.
As previously discussed in Sec. 4.2, the omni-modal instruction data is formulated as $D_{\text{omni}}=\{(x_{V},u_{S},x^{o}_{T},c^{o}_{\text{style}},u^{o}_{S})_{i}\}_{i=1}^{N}$. In details, the textual outputs $x_{T}^{o}=(x_{T}^{o^{1}},x_{T}^{o^{2}})$ contain the transcribed textual instructions $x_{T}^{o^{1}}$ and the textual responses $x_{T}^{o^{2}}$. The styles labels $c^{o}_{\text{style}}=(c^{o}_{\text{emo}},c^{o}_{\text{p}})$ include the emotion and pitch labels, respectively.
Mathematical formulation of chain of modality.
Based on the notations above, the sequential chain of modality approach can be mathematically formulated by decomposing the conditional likelihood of the desired outputs $(x_{T}^{o^{1}},x_{T}^{o^{2}},c^{o}_{\text{emo}},c^{o}_{\text{p}},u^{o}_{S})$ given the inputs $(x_{V},u_{S})$. Specifically, let $z_{1}=x_{T}^{o^{1}},z_{2}=x_{T}^{o^{2}},z_{3}=c^{o}_{\text{emo}},z_{4}=c^{o}_{\text{p}},\text{and},z_{5}=u^{o}_{S}$, the decomposition is expressed as:
$\mathbb{P}(x_{T}^{o^{1}},x_{T}^{o^{2}},c^{o}_{\text{emo}},c^{o}_{\text{p}},u^{o}_{S}\mid x_{V},u_{S})=\prod_{i=1}^{5}\mathbb{P}(z_{i}\mid z_{1:i-1},x_{V},u_{S}).$ (3)
Dataset Source # Examples Emotions Pitches
Neutral Happy Sad Angry Normal Low High
Speech-Image
-EN(train)
ALLAVA 71,474 58,506 12,412 516 40 70,962 392 120
Speech-Image
-EN(test)
ALLAVA 1,056 434 274 300 48 44 176 16
Speech-Image
-ZH(train)
ALLAVA
38,260 29,893 7,680 607 80 36,363 624 1,273
Speech-Image
-ZH(test)
ALLAVA
616 96 193 190 137 381 177 58
Speech-Text
-EN(train)
ShareGPT 73,658 42,334 20,946 4,674 5,704 60,352 5,518 7,788
Speech-Text
-EN(test)
ShareGPT 1,400 200 400 400 400 582 422 422
Speech-Text
-ZH(train)
In-house 47,852 29,769 16,405 1,362 316 42,356 4,379 1,117
Speech-Text
-ZH(test)
In-house 684 96 196 196 196 458 134 92
Table 4: Statistics of the EMOVA speech instruction tuning datasets.
Settings Stage-1 Stage-2 Stage-3
Vision Resolution Min $56^{2}$, Max $1792^{2}$ Min $56^{2}$, Max $1792^{2}$ Min $56^{2}$, Max $1792^{2}$
# Tokens Min $4$, Max $4096$ Min $4$, Max $4096$ Min $4$, Max $4096$
Data Dataset LCS EMOVA-Alignment (Fig. 9) EMOVA-SFT (Fig. 8)
# Samples 558K 7.4M 4.4M
Training Trainable Projector Full Model (Half ViT) Full Model
Batch Size 256 256 128
Context Length 8192 8192 8192
LR: $p(\cdot)$ $1\times 10^{-3}$ $2\times 10^{-5}$ $2\times 10^{-5}$
LR: $v(\cdot)$ $2\times 10^{-5}$ $2\times 10^{-6}$
LR: $f(\cdot)$ $2\times 10^{-5}$ $2\times 10^{-5}$
Epoch 1 1 1
Table 5: Detailed configuration for different training stages of EMOVA. The table illustrates the vision configurations, dataset characteristics, and training hyperparameters.
Appendix C More on Benchmark Evaluation
To thoroughly evaluate our model’s vision-language abilities, 15 benchmarks covering four different aspects of the real-life scenarios are utilized for a comprehensive assessment across multiple domains. Moreover, Automatic Speech Recognition (ASR) and Text-to-speech (TTS) are adopted to evaluate speech-language abilities.
Document/chart understanding and OCR abilities.
Benchmarks including TextVQA [76], ChartQA [61], DocVQA [62], InfoVQA [63], and OCRBench [51], focus on recognition and understanding of structured data (e.g., charts, documents, and characters), challenging the model to extract, comprehend, and reason with structural and textual data. We adopt the corresponding test set for evaluation.
General image perception and understanding.
MME [18], MMBench [50], SEED-Bench-Image [39], MM-Vet [90], and RealWorldQA [86] evaluate models on perception and reasoning among general vision domains, providing a comprehensive measurement of models’ real-world generalization abilities.
Mathematical problem solving.
MathVista (testmini) [57] and MathVerse (vision-mini) [96] evaluate the model’s ability to deal with diverse mathematical problems, including both arithmetic and complex reasoning questions across different levels of complexity.
Science understanding.
MMMU [91], ScienceQA-Img [56] and AI2D [33] are used to assess models’ ability to deal with scientific questions and diagrams, which requires reasoning across various subjects and interpreting structured visual content.
Automatic speech recognition (ASR).
We utilize the test-clean set of LibriSpeech [72] for English, reporting the Word Error Rate (WER) as the evaluation metric. For Chinese, evaluation is conducted on the test set of AISHELL-2 [15], using the Character Error Rate (CER). Both WER and CER assess ASR performance, calculated by comparing the recognized texts with the ground-truth transcripts.
Text-to-speech (TTS).
To evaluate the TTS abilities, we first prompt EMOVA to generate speech units, which are then converted to the speech waveforms by our U2S detokenizer. Using the synthesized speech as input, we conduct ASR with Whisper-large-v3 and Paraformer-zh for English and Chinese, respectively, to obtain transcribed texts. We then compute the WER and CER between the ground truth texts and the transcribed texts as metrics for TTS. The resulting metrics are denoted as TTS-WER and TTS-CER for English and Chinese.
Appendix D More on Evaluation of Speech-Language Capabilities
D.1 Calculation of Evaluation Metrics
End-to-end spoken dialogue score.
We prompt GPT-4o with the original question $x_{T}^{o^{1}}$, the ground-truth text answer $x_{T}^{o^{2}}$ and the transcribed text from the generated speech, to obtain a score ranging from 0 to 10 and report an average of them. The prompt can be found in Fig. 25.
Unit-input-text-output score.
Similar to end-to-end spoken dialogue score, but we use the predicted text response $\tilde{x_{T}^{o^{2}}}$ as answer instead of the transcribed text from the generated speech, to obtain a score ranging from 0 to 10 and report an average of them. See the prompt in Fig. 24.
Text-input-text-output score.
The prompt can be found in Fig. 24.
Style label classification accuracy.
We use GPT-4o to decide whether the style predictions $\tilde{c_{\text{emo}}^{o}},\tilde{c_{\text{p}}^{o}}$ are correct given the transcribed instruction $\tilde{x_{T}^{o^{1}}}$ and the predicted text response $\tilde{x_{T}^{o^{2}}}$. The prompt can be found in Fig. 26.
Emotion controllablity
of our U2S detokenizer is assessed by providing texts to LLM to generate corresponding units (i.e., TTS), which, along with the given conditional emotion labels, are then fed into our U2S detokenizer to synthesize speech. We choose the female voice due to its large variation of styles. We evaluate on 4 commonly-seen emotions, i.e.,“neutral”, “happy”, “sad”, and “angry”. We synthesize 200 speech utterances for testing, with 50 utterances per emotion. The output speeches are analyzed by a Speech Emotion Recognition (SER) model named Emotion2Vec [59], which identifies the emotion with the greatest likelihood among these four emotions.
[…]
Table 6: Comparison on the ASR and TTS benchmarks.
D.2 Comparison with other Omni Models
Experimental results of ASR and TTS are reported in Table 2 and 6. EMOVA achieves significant improvements over other omni-modal models (i.e., AnyGPT and VITA), even surpassing its SLLM counterpart Mini-Omni [87], demonstrating the effectiveness of semantic-acoustic disentanglement and omni-modal mutual benefits. For the first time, our EMOVA obtains state-of-the-art performance on both the vision-language and speech benchmarks simultaneously.
Appendix E More on Vision-language Architecture
ViT ViT LR Template MME MMBench SEED-Image TextVQA ChartQA DocVQA InfoVQA OCRBench ScienceQA-Img AI2D
Full 2×$\times 10^{-6}$ QA 1928 68.8 72.5 64.3 29.9 45.2 28.7 495 76.3 61.8
Half 2×$\times 10^{-6}$ QA 1838 71.3 72.8 63.3 31.4 46.0 28.5 489 76.1 63.7
Frozen 2×$\times 10^{-6}$ QA 1887 68.8 72.2 61.3 30.2 44.7 28.0 478 75.9 62.8
Half 2×$\times 10^{-5}$ QA 1833 68.3 73.1 62.2 27.8 42.4 27.3 461 74.7 63.4
Half 2×$\times 10^{-6}$ Plain 1909 70.1 72.0 61.5 24.5 38.9 30.1 410 77.0 63.6
Table 7: Ablation on the ViT configurations and templates for vision-language alignment.
This section analyzes the pre-training configurations of the vision encoder and the prompt template during Stage 2, as shown in Table. 7. Our final selection is the colored setting. We find that training the ViT model with half of the deeper parameters [9] with a learning rate of 2×$\times 10^{-6}$ [48] yields the best performance. Furthermore, we compare the QA template with the plain template [49] and find that the QA template is superior for pre-training.
Appendix F Limitations
Duplex modeling.
In the current version, EMOVA can only process either visual/speech/text inputs or produce speech/text outputs at the same time. For a communication experience that mirrors human interaction, handling inputs and outputs simultaneously is crucial. Recent works like VITA [19] and LSLM [60] have begun to explore duplex modeling. VITA focuses on recognizing speech in noisy environments during the generation process to facilitate timely responses. LSLM, on the other hand, attempts to halt speech production when it detects a command or voice. Recently, a ground-breaking work named Moshi [13] develops a model supporting fully duplex modeling. The adeptness at simultaneously managing the information streams from both the user and the assistant allows Moshi to converse with human beings in real-time scenarios.
However, incorporating emotions into this duplex modeling presents additional challenges. Emotional states can fluctuate throughout a conversation, and how to generate appropriate responses given the user’s previous and current emotional cues has not been thoroughly investigated. We will dive into this topic in the future work.
Direct unit-to-unit generation.
Although the speech units have served as the speech representation, they are predominantly adopted in conjunction with text-based assistance [95]. However, the direct generation from unit to unit without text assistance is an area that has not been extensively explored. In [37], speeches from the source language are directly translated into speech units of the target language for speech-to-speech translation. Similarly, [68] builds a language model directly on speech units, enabling spoken dialogue generation from raw audio. Both works develop models in speech-only data.
In the current version of EMOVA, the text modality is integrated into the speech generation process to transfer textual knowledge to the speech modality, thereby enhancing the correctness of speech responses. In the future, we will strengthen the model’s direct unit-to-unit generation capabilities to boost the speed of speech generation and augment the model’s comprehension of speech modality.
Vision configurations.
Currently, we only utilize a single vision encoder pre-trained via a vision-language manner, while recent works have shown effectiveness by combining vision encoders pre-trained by different manners (e.g., self-supervised pre-training [52]) and architectures (e.g., MoE [98, 54]). We prioritize visual understanding in this work, while the incorporation of (controllable) visual generation [8, 22, 42, 83, 53, 23, 24] is also appealing to better empower EMOVA in real-life applications [44, 41, 29, 85]. Digging into its robustness towards noisy vision inputs [28] is also an appealing direction.
Comparison with Emu3 [82].
Both our EMOVA and Emu3 build end-to-end MLLMs with discrete tokenization. Emu3 discretizes visual data, enabling visual understanding and generation, while our EMOVA discretizes speech data with a continuous visual encoder, building an Omni-modal LLM with visual, text, and speech abilities simultaneously.
Appendix G Qualitative Results
Advanced Vision-language Abilities.
Our EMOVA exhibits advanced vision-language abilities in humor understanding (Fig. 10), numerical calculations (Fig. 11), coding (Fig. 12), geometry problem solving (Fig. 16).
(Omni-modal) Emotional Spoken Dialogue.
EMOVA can engage in omni-modal emotional spoken dialogue (Figs. 5, 13, 14, 18). For demonstration purposes, we present only the speech transcriptions [^2] For speech files, please refer to our project page. of the instruction and EMOVA’s response. The text highlighted in blue indicates the style labels predicted by EMOVA, reflecting the emotion and pitch of generated speeches.
Omni-modal Spoken Dialogue with Structural Data Understanding.
EMOVA can perform structural data understanding even in spoken dialogue. The data types include PDF files (Fig. 15), movie posters (Figs. 17 and 19), personal résumés (Fig. 20), charts (Fig. 21), and websites (Fig. 22). Remarkably, EMOVA is not specifically trained on spoken dialogue data involving these data types. This suggests that our chain of modality data introduced in Sec. 4.2 effectively activates both vision-language abilities and speech understanding acquired during the text-centric alignment stage.
Figure 10: EMOVA exhibits advanced vision-language abilities in humor sense.
Figure 11: EMOVA exhibits advanced vision-language abilities in numerical calculations.
Figure 12: EMOVA exhibits advanced vision-language abilities in coding.
Figure 13: EMOVA engages in spoken dialogue with a cheerful tone.
Figure 14: EMOVA engages in emotional spoken dialogue.
Figure 15: EMOVA engages in omni-modal spoken dialogue with structural data understanding.
Figure 16: EMOVA exhibits advanced vision-language abilities in math.
Figure 17: EMOVA engages in omni-modal spoken dialogue with structural data understanding (i.e., poster).
Figure 18: EMOVA engages in omni-modal emotional spoken dialogue.
Figure 19: EMOVA engages in omni-modal spoken dialogue with structural data understanding (i.e., movie posters).
Figure 20: EMOVA engages in omni-modal spoken dialogue with structural data understanding (i.e., résumé).
Figure 21: EMOVA engages in omni-modal spoken dialogue with structural data understanding (i.e., chart).
Figure 22: EMOVA engages in omni-modal spoken dialogue with structural data understanding (i.e., website).
Figure 23: Prompt used to obtain style labels of the speech instruction dataset.
Figure 24: Prompt used to obtain Unit-Input-Text-Output Score and Text-Input-Text-Output Score.
Figure 25: Prompt used to obtain End-to-end Spoken Dialogue Score.
Figure 26: Prompt used to obtain Classification Accuracy of Style Label.